OpenStudy (anonymous):

A turntable with a moment of inertia of 0.2 kgm^2 and a radius of 0.3 meters starts from rest. After 6 seconds, it's angular speed is 30 degrees radians/second. A. What is the angular acceleration? B. What angle does it turn through during the six second interval? C. What is the torque on the turntable?

OpenStudy (michele_laino):

hint: anglular acceleration \alpha is: \[\alpha = \frac{\omega }{{\Delta t}}\] where \omega is the gained angular speed, and \Delata t is time needed to gain that angular speed

OpenStudy (michele_laino):

Part B: the requested angle \theta, is given by the subsequent formula: \[\theta = \frac{{\omega {{\left( {\Delta t} \right)}^2}}}{2}\]

OpenStudy (michele_laino):

Part C: whereas the requested torque M, is: \[M = I\omega \] where I is the moment of inertia